# IC/DC: Surpassing Heuristic Solvers in Combinatorial Optimization with Diffusion Models

### v1

@@ 4 Architecture
To effectively encode the problem represented by these matrices, we adopt the dual graph attentional layer structure from MatNet (Kwon et al. 2021), but replace the attention layer with a modified version of graph attention networks (GAT, Veličković et al. 2017) that is specifically designed to process a bipartite graph. The problem encoder consists of $L$ layers, where each layer takes $(A,B,D)$ as input and outputs updated features $(A^{\prime},B^{\prime})$. The outputs of the final layer are then passed to the denoiser. The bottom-right side of Figure 2 illustrates this problem encoder. Detailed equations of problem encoder layer are provided in Appendix B.2.
@@
Building on recent empirical successes (Joshi et al. 2020; Qiu, Sun, and Yang 2022), we extend the anisotropic graph neural network (AGNN) to handle bipartite graphs, allowing us to consider two distinct sets of items, and use it as the denoiser. The input embedding layer maps each element of the noisy solution matrix $X_{t}$ and the timestep $t$ into $d$-dimensional features. These embeddings are then passed to the AGNN, along with the problem embedding $A^{\prime}$ and $B^{\prime}$. After $L^{\prime}$ layers of AGNN, the embedded solution matrix with updated features is passed through a linear layer to produce the denoised solution matrix $X_{0}$. The bottom-left side of Figure 2 illustrates this denoiser. Detailed equations of denoiser layer are provided in Appendix B.3.
@@ 5 Experiments
Makespan $\downarrow$ Gap $\downarrow$ Time $\downarrow$ Makespan $\downarrow$ Gap $\downarrow$ Time $\downarrow$
Baseline CP-SAT (OR-tool) 28.110 0% (1m) 65.900 0% (3m)
(Meta-)heuristics Random 59.468 71.612% (1m) 138.683 71.153% (3m)
Shortest Job First 34.250 19.692% (1m) 72.986 10.204% (2m)
Genetic Algorithm 30.211 7.205% (2h) 68.912 4.468% (4h)
Particle Swarm Opt. 29.794 5.817% (1h) 69.133 4.788% (3h)
Auto-regressive MatNet 29.588 5.123% (0s) 67.092 1.793% (1s)
MatNet (×$\times 8$) 28.491 1.346% (1s) 66.129 0.347% (3s)
MatNet (×$\times 32$) 28.353 0.860% (2s) 66.105 0.311% (16s)
MatNet (×$\times 128$) 28.283 0.615% (15s) 66.020 0.182% (1m)
[…]
Ours IC/DC 29.168 3.694% (7s) 66.677 1.172% (10s)
IC/DC (×$\times 8$) 28.300 0.674% (15s) 66.196 0.448% (28s)
IC/DC (×$\times 32$) 28.190 0.284% (46s) 66.055 0.235% (1m)
IC/DC (×$\times 128$) 28.150 0.142% (3m) 65.974 0.112% (5m)
Table 1: (Results for PMSP-20 and PMSP-50) The top row (labeled as baseline) represents the results from an off-the-shelf solver and is used as a reference for calculating the performance gap. $\infty$ indicates that the generated solutions are infeasible. Results where our model outperforms the previous SOTA learning-based method are highlighted in bold.
@@
[…]
Ours IC/DC 1.609 4.802% (1s) 1.614 3.969% (15s)
IC/DC (×$\times 8$) 1.549 0.971% (8s) 1.566 0.928% (2m)
IC/DC (×$\times 32$) 1.534 -0.026% (31s) 1.547 -0.235% (8m)
IC/DC (×$\times 128$) 1.530 -0.235% (2m) 1.543 -0.532% (30m)
Table 2: (Results for ATSP-20 and ATSP-50) The top row (labeled as baseline) represents the results from an off-the-shelf solver and is used as a reference for calculating the performance gap. $\infty$ indicates that the generated solutions are infeasible.
@@ 5.1 Demonstrated Problems
$\text{score}_{\text{PMSP}}(X_{0},c)=-\max_{m}\sum_{j}[X_{0}\odot P]_{j,m}.$
@@
$\text{score}_{\text{ATSP}}(X_{0},c)=-\sum_{i,j}{[X_{0}\odot D]_{i,j}}.$
The solution matrix that is not a Hamiltonian cycle is considered infeasible. In accordance with Kwon et al. (2021) we employ tmat-class ATSP instances (see Appendix D.1).
@@ 5.2 Main Results
We evaluated the baselines and the proposed algorithm using two Intel Xeon Gold 6330 CPUs and an RTX 3090 GPU for both PMSP and ATSP. For the evaluation, 1000 problem instances were randomly generated using a standard generation process (see appendix D.1). To fully leverage the stochastic nature of generative learning-based methods, we also evaluated these methods by generating multiple samples (×$\times n$) for each problem instance and selecting the one with the best score. For MatNet (Kwon et al. 2021), we followed the authors’ implementation, including instance augmentation, which yielded better result. As a problem-specific search process has not been studied on PMSP and ATSP, for diffusion-based methods, we report simple discrete diffusion models, either trained with supervised learning (Sun and Yang 2023) or with reinforcement learning (Black et al. 2023). For further experimental details, please refer to Appendix C and Appendix D.
@@
We evaluated our method on PMSP-20 and PMSP-50, where the numbers 20 and 50 correspond to the number of jobs in each instance, with the number of machines fixed at 4. Our approach is compared against various baselines, including (meta-)heuristics, auto-regressive, and diffusion-based methods. Details of these baselines can be found in Appendix C.2.
As shown in Table 1, IC/DC achieves the smallest optimality gap among learning-based methods. IC/DC demonstrates a $0.142\%$ performance gap compared to CP-SAT, whereas the previous SOTA, MatNet, shows a $0.615\%$ gap on PMSP-20. On PSMP-50 IC/DC reduces the gap from MaNet’s $0.182\%$ to $0.112\%$. Although IC/DC has a slower inference speed compared to auto-regressive methods, its high-quality solutions remain highly competitive with other baselines.
@@
We evaluated our method on ATSP-20 and ATSP-50, where the number 20 and 50 correspond to the number of cities. IC/DC generates solutions that not only surpass those generated by SOTA learning-based methods but also outperform those obtained through (meta-)heuristics and off-the-shelf solvers. IC/DC achieves a $-0.235\%$ gap on ATSP-20 and a $-0.532\%$ gap on ATSP-50, demonstrating a negative gap compared to a powerful CPLEX solver—which no other learning-based method has achieved. This substantial improvement positions IC/DC as a new SOTA method for addressing these less-studied CO problems, highlighting its potential.
5.3 Discussion
Comparison with auto-regressive method
As discussed earlier, the performance of auto-regressive models is limited by their sequential decision-making process, which locks in previously made decisions, preventing any revisions. While IC/DC employs a similar auto-regressive approach for feasibility-enforced generation, it also benefits from the accuracy of diffusion-based methods by iteratively refining the solution through $T-1$ denoising steps before the final auto-regressive generation. Moreover, this two-stage generation process enables IC/DC to achieve much greater sample diversity compared to auto-regressive models, as the denoising process provides varied foundations for solutions before the auto-regressive generation.
The results that support these arguments can be found in both tables, particularly in Table 2. While IC/DC’s performance lags behind MatNet when using a single generation to solve the problem, its quickly surpasses MatNet as the number of samples increases, highlighting the high diversity of IC/DC’s samples. Upon closer examination, we observed that the stochasticity of MatNet primarily stems from the starting point of the auto-regressive generation (i.e., the selection of the initial city to begin the tour). In contrast, IC/DC’s denoising process offers diverse backbones for the auto-regressive generation, resulting in varied samples even when starting from the same initial city.
Comparison with diffusion-based method
Previously proposed diffusion-based methods (Qiu, Sun, and Yang 2022; Sun and Yang 2023; Min, Bai, and Gomes 2024) have employed problem-specific search processes to enforce feasibility and enhance performance. However, developing such search processes with strong performance and feasibility guarantees typically requires significant expertise in the specific CO problem being addressed. As the complexity of CO problem increases, such as with ATSP or more complex TSP variants compared to TSP, the cost of implementing these search algorithms becomes even greater.
@@ 5.3 Discussion
As shown in Table 1 and Table 2, simple implementations of diffusion models, whether trained with supervised learning (Sun and Yang 2023) or reinforcement learning (Black et al. 2023), struggle to generate feasible solutions without the aid of problem-specific search processes, which have not been studied for these particular CO problems. This highlights the critical importance of our feasibility-enforced generation process, which integrates the flexibility of auto-regressive methods into diffusion-based methods, allowing for the straightforward imposition of various constraints. In this regard, the proposed IC/DC approach not only demonstrates superior performance compared to previous learning-based methods, but also significantly broadens the applicability of diffusion-based methods to a wide range of less-explored CO problems with diverse constraints.
@@ 6 Conclusion
We propose a diffusion-based approach named IC/DC to tackle the CO problems involving two distinct sets of items. Our algorithm demonstrates superior performance in both PMSP and ATSP, outperforming existing baselines and successfully generating feasible solutions that cannot be easily achieved with simple diffusion-based methods. We emphasize that this is the first study to successfully adapt a diffusion-based method for CO problems involving two distinct sets of items. We believe the proposed IC/DC framework has significant potential for generlizability. For example, in Appendix D.4, we demonstrate its capability to address real-world CO problems with sophisticated item and relationship features.
Despite its strong performance, IC/DC has a notable limitation: the GAT-based encoder we use requires $O(\max(|\mathcal{A}|,|\mathcal{B}|)^{2})$ memory complexity. This demands substantial memory resources when training on large instances. To address these challenges, we plan to explore memory-efficient techniques in future work, such as those introduced by Zhu et al. (2024).
@@ A.1 Joint Variational Upper Bound with EUBO
We derive the upper bound using the Evidence Upper Bound (EUBO) (Ji and Shen 2019)
@@
$D_{KL}(\tilde{q}(X|c)\|p_{\theta}(X|c))$ $=\int\tilde{q}(X|c)\log\frac{\tilde{q}(X|c)}{p_{\theta}(X|c)}dX$
@@ A.2 KL divergence to Reward-weighted Diffusion Loss
We derive the reward-weighted diffusion loss from the objective in CLONING step. We aim to minimize the Kullback-Leibler (KL) divergence $D_{KL}(\tilde{q}(X|c)\parallel p_{\theta}(X|c))$ between the surrogate target and the distribution $p$ parameterized by $\theta$. however since evaluating the log-likelihood of the generative model $p_{\theta}(X|c)$ is difficult (Kingma and Welling 2013), we instead utilize a joint variational upper bound (see App. A.1):
$\mathrm{D}_{\mathrm{KL}}(\tilde{q}(X|c)\parallel p_{\theta}(X|c))\leq\mathrm{D}_{\mathrm{KL}}(\tilde{q}(X,Z|c)\parallel p_{\theta}(X,Z|c))$
@@
Therefore, we focus on minimizing the right-hand side of eq (A.2). By applying the diffusion process where $Z=X_{1:T}$, $X=X_{0}$, and $T$ is the diffusion step, the forward process is defined by $q(X_{1:T}|X_{0})$ with $X_{0}\sim\tilde{q}(X_{0}|c)$. We minimize the Reward-weighted Diffusion Loss $\mathcal{L}_{\text{RWD}}$:
$\mathcal{L}_{\text{RWD}}$ $=D_{KL}(\tilde{q}(X_{0:T}|c)\parallel p_{\theta}(X_{0:T}|c))$
$=D_{KL}(\tilde{q}(X_{0},X_{1:T}|c)\parallel p_{\theta}(X_{0:T}|c))$
$=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{1:T}\sim q(X_{1:T}|X_{0})}\left[\log\frac{q(X_{1:T}|X_{0})\tilde{q}(X_{0}|c)}{p_{\theta}(X_{0:T}|c)}\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[\log\frac{q(X_{1:T}|X_{0})}{p_{\theta}(X_{0:T}|c)}+\log\tilde{q}(X_{0}|c)\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[\log\frac{\prod_{t=1}^{T}q(X_{t}|X_{t-1})}{p_{\theta}(X_{T}|c)\prod_{t=1}^{T}p_{\theta}(X_{t-1}|X_{t},c)}+\log\tilde{q}(X_{0}|c)\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[-\log p_{\theta}(X_{T}|c)+\sum_{t=1}^{T}\log\left(\frac{q(X_{t}|X_{t-1})}{p_{\theta}(X_{t-1}|X_{t},c)}\right)+\log\tilde{q}(X_{0}|c)\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[-\log p_{\theta}(X_{T}|c)+\sum_{t=2}^{T}\log\left(\frac{q(X_{t}|X_{t-1})}{p_{\theta}(X_{t-1}|X_{t},c)}\right)+\log\left(\frac{q(X_{1}|X_{0})}{p_{\theta}(X_{0}|X_{1},c)}\right)+\log\tilde{q}(X_{0}|c)\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\Bigg{[}-\log p_{\theta}(X_{T}|c)$
$\quad\quad\quad\quad\quad\quad\quad\ \ +\sum_{t=2}^{T}\log\left(\frac{q(X_{t-1}|X_{t},X_{0})}{p_{\theta}(X_{t-1}|X_{t},c)}\cdot\frac{q(X_{t}|X_{0})}{q(X_{t-1}|X_{0})}\right)+\log\left(\frac{q(X_{1}|X_{0})}{p_{\theta}(X_{0}|X_{1},c)}\right)+\log\tilde{q}(X_{0}|c)\Bigg{]}$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[\log\frac{q(X_{T}|X_{0})}{p_{\theta}(X_{T}|c)}+\sum_{t=2}^{T}\log\left(\frac{q(X_{t-1}|X_{t},X_{0})}{p_{\theta}(X_{t-1}|X_{t},c)}\right)-\log p_{\theta}(X_{0}|X_{1},c)+\log\tilde{q}(X_{0}|c)\right]$
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\Bigg{[}\underbrace{D_{KL}(q(X_{T}|X_{0})\|p_{\theta}(X_{T}|c))}_{L_{T}}+\sum_{t=2}^{T}\underbrace{D_{KL}(q(X_{t-1}|X_{t},X_{0})\|p_{\theta}(X_{t-1}|X_{t},c))}_{L_{t-1}}$
$\quad\quad\quad\quad\quad\quad\quad\ \ +\underbrace{D_{KL}(\tilde{q}(X_{0}|c)\|p_{\theta}(X_{0}|X_{1},c))}_{L_{0}}\Bigg{]}$
If $T$ is sufficiently large, $L_{T}$ will approach zero (Austin et al. 2021). Also, $L_{0}$ can be derived as CE loss:
$L_{0}$ $=\sum\tilde{q}(X_{0}|c)\log\frac{\tilde{q}(X_{0}|c)}{p_{\theta}(X_{0}|X_{1},c)}=\sum\tilde{q}(X_{0}|c)\log\tilde{q}(X_{0}|c)-\log p_{\theta}(X_{0}|X_{1},c))$
$=CE(\tilde{q},p_{\theta})-H(\tilde{q})\approx\mathcal{L}_{\text{prd}}-H(\tilde{q})$
@@
$\therefore\mathcal{L}_{\text{RWD}}\approx\mathcal{L}_{\text{VB}}+\mathcal{L}_{\text{prd}}$

### v4

@@ 4 Architecture
To effectively encode the problem represented by these matrices, we adopt the dual graph attentional layer structure from MatNet [19], but replace the attention layer with a modified version of graph attention networks (GAT, [28] 2017) that is specifically designed to process a bipartite graph. The problem encoder consists of $L$ layers, where each layer takes $(A,B,D)$ as input and outputs updated features $(A^{\prime},B^{\prime})$. The outputs of the final layer are then passed to the denoiser. The bottom-right side of Fig. 2 illustrates this problem encoder. Detailed equations of problem encoder layer are provided in Sec. B.2.
@@
Building on recent empirical successes [15, 22], we extend the anisotropic graph neural network (AGNN) to handle bipartite graphs, allowing us to consider two distinct sets of items, and use it as the denoiser. The input embedding layer maps each element of the noisy solution matrix $X_{t}$ and the timestep $t$ into $d$-dimensional features. These embeddings are then passed to the AGNN, along with the problem embedding $A^{\prime}$ and $B^{\prime}$. After $L^{\prime}$ layers of AGNN, the embedded solution matrix with updated features is passed through a linear layer to produce the denoised solution matrix $X_{0}$. The bottom-left side of Fig. 2 illustrates this denoiser. Detailed equations of denoiser layer are provided in Sec. B.3.
@@ 5 Experiments
3 $\times$ 20 4 $\times$ 20 5 $\times$ 20 3 $\times$ 50 4 $\times$ 50 5 $\times$ 50
Method M.S. $\downarrow$ Gap $\downarrow$ Time $\downarrow$ M.S. $\downarrow$ Gap $\downarrow$ Time $\downarrow$ M.S. $\downarrow$ Gap $\downarrow$ Time $\downarrow$ M.S. $\downarrow$ Gap $\downarrow$ Time $\downarrow$ M.S. $\downarrow$ Gap $\downarrow$ Time $\downarrow$ M.S. $\downarrow$ Gap $\downarrow$ Time $\downarrow$
Baseline CP-SAT 42.63 0% (1m) 28.11 0% (1m) 20.58 0% (3m) 102.06 0% (2m) 65.90 0% (3m) 47.03 0% (5m)
(Meta-)heuristics SJF 48.03 11.91% (1m) 34.25 19.70% (1m) 27.22 27.77% (1m) 108.44 6.06% (2m) 72.99 10.20% (2m) 54.49 14.69% (3m)
GA 44.21 3.64% (2h) 30.21 7.21% (2h) 23.04 11.28% (2h) 104.51 2.37% (4h) 68.91 2.47% (4h) 50.48 7.08% (4h)
PSO 43.84 2.80% (1.5h) 29.79 5.82% (1.5h) 22.70 9.79% (1.5h) 104.70 2.55% (3h) 69.13 4.79% (3h) 50.71 7.53% (3h)
Autoregressive MatNet 43.78 2.67% (0s) 29.59 5.12% (0s) 21.36 3.70% (1s) 103.16 1.07% (2s) 67.09 1.79% (1s) 47.67 1.36% (6s)
MatNet ($\times$8) 42.87 0.56% (1s) 28.49 1.35% (1s) 20.87 1.41% (2s) 102.42 0.34% (5s) 66.13 0.35% (3s) 47.31 0.59% (17s)
MatNet ($\times$32) 42.73 0.24% (2s) 28.35 0.86% (2s) 20.81 1.12% (6s) 102.23 0.16% (9s) 66.11 0.31% (16s) 47.19 0.33% (39s)
MatNet ($\times$128) 42.67 0.10% (9s) 28.28 0.62% (15s) 20.68 0.50% (37s) 102.14 0.08% (42s) 66.02 0.18% (1m) 47.12 0.18% (5m)
Heatmap-based SL ($\times$128) 45.31 6.09% (1m) 31.12 10.16% (2m) 23.76 14.34% (2m) 107.88 5.54% (3m) 71.95 8.78% (3m) 53.65 13.15% (4m)
RL ($\times$128) 63.94 39.997% (9m) 50.93 57.74% (9m) 43.39 71.323% (10m) 152.86 39.858% (22m) 120.08 58.26% (24m) 107.719 78.435% (28m)
Ours IC/DC 43.06 1.00% (3s) 29.17 3.69% (7s) 21.16 2.80% (4s) 102.87 0.79% (14s) 66.68 1.17% (10s) 48.18 2.41% (15s)
IC/DC ($\times$8) 42.67 0.16% (5s) 28.30 0.67% (15s) 20.75 0.80% (7s) 102.27 0.21% (20s) 66.20 0.45% (28s) 47.44 0.87% (45s)
IC/DC ($\times$32) 42.64 0.02% (21s) 28.19 0.28% (46s) 20.66 0.39% (25s) 102.13 0.06% (50s) 66.06 0.24% (1m) 47.22 0.39% (2m)
IC/DC ($\times$128) 42.63 0% (47s) 28.15 0.14% (3m) 20.61 0.15% (2m) 102.08 0.01% (3m) 65.97 0.11% (5m) 47.11 0.17% (7m)
Table 1: (Results for PMSP-20 and PMSP-50) The table shows Makespan (M.S.), Gap, and Time for 3 × 20, 4 × 20, and 5 × 20, along with PMSP-50 (3 × 50, 4 × 50, 5 × 50).
@@ 3 $\times$ 20 4 $\times$ 20 5 $\times$ 20 3 $\times$ 50 4 $\times$ 50 5 $\times$ 50
Heatmap-based Diffusion (SL) (×$\times 128$) $1.599$ $4.179\%$ (5m) $1.684$ $8.215\%$ (30m)
Diffusion (RL) (×$\times 128$) 3.331 $73.875\%$ (6m) 4.334 $94.579\%$ (32m)
Diffusion (SL+RL) (×$\times 128$) 1.589 $3.508\%$ (5m) 1.679 $7.964\%$ (30m)
Ours IC/DC 1.609 4.802% (1s) 1.619 4.266% (15s)
IC/DC (×$\times 8$) 1.548 0.909% (8s) 1.570 1.243% (2m)
IC/DC (×$\times 32$) 1.546 0.779% (31s) 1.564 0.828% (8m)
IC/DC (×$\times 128$) 1.534 0% (2m) 1.553 0.113% (20m)
Table 2: (Results for ATSP-20 and ATSP-50) The top row (labeled as baseline) represents the results from an off-the-shelf solver and is used as a reference for calculating the performance gap.
@@ 5.1 Demonstrated Problems
$\text{score}_{\text{PMSP}}(X_{0},c)=-\max_{m}\sum_{j}[X_{0}\odot P]_{j,m}.$ (18)
@@
$\text{score}_{\text{ATSP}}(X_{0},c)=-\sum_{i,j}{[X_{0}\odot D]_{i,j}}.$ (19)
The solution matrix that is not a Hamiltonian cycle is considered infeasible. In accordance with Kwon et al. [19] we employ tmat-class ATSP instances (see Sec. D.1).
Figure 3: (Left) shows the optimality gap among different methods in ATSP. Each of the methods used a unified sampling size of 128. The * indicates the training distribution (20 nodes). (Center) shows the optimality gap across sampling step size in PMSP (3 machines and 20 jobs). (Right) shows the inference time across sampling step size.
@@ 5.2 Main Results
For the evaluation, 1000 problem instances were randomly generated using a standard generation process (see Sec. D.1). To fully leverage the stochastic nature of generative learning-based methods, we also evaluated these methods by generating multiple samples (×$\times n$) for each problem instance and selecting the one with the best score. For MatNet [19], we followed the authors’ implementation, including instance augmentation, which yielded better result. As a problem-specific search process has not been studied on PMSP and ATSP, for heatmap-based methods, we report simple discrete diffusion models, either trained with supervised learning [26] or with reinforcement learning [5]. For further experimental details, please refer to Appendix C and Appendix D.
@@
We evaluated our method on PMSP-20 and PMSP-50, where the numbers 20 and 50 correspond to the number of jobs in each instance, with the number of machines fixed at 4. Our approach is compared against various baselines, including (meta-)heuristics, autoregressive, and heatmap-based methods. Details of these baselines can be found in Sec. C.2. As shown in Tab. 1, IC/DC achieves the smallest optimality gap among learning-based methods. IC/DC demonstrates a $0.142\%$ performance gap compared to CP-SAT, whereas the previous SOTA, MatNet, shows a $0.615\%$ gap on PMSP-20. On PSMP-50 IC/DC reduces the gap from MatNet’s $0.182\%$ to $0.112\%$.
With combination of feasibility-enforced generation and heatmap-based iterative refinement, ID/DC achieves greater sample diversity, as the denoising process provides varied foundations for solutions before the autoregressive generation. While IC/DC’s performance lags behind MatNet when using a single generation to solve the problem, it quickly surpasses MatNet as the number of samples increases (Tab. 1). Upon closer examination, we observed that the stochasticity of MatNet primarily stems from the starting point of the autoregressive generation (i.e., the assignment of the initial job for scheduling). In contrast, IC/DC’s denoising process offers diverse backbones for the autoregressive generation, resulting in varied samples even when starting from the same initial assignment. Although IC/DC has a slower inference speed compared to autoregressive methods, its high-quality solutions remain highly competitive with other baselines.
@@
We evaluated our method on ATSP-20 and ATSP-50, where the number 20 and 50 correspond to the number of cities. To our best knowledge, IC/DC is the first framework among NCO to achieve optimality gap with $0\%$ in ATSP-20. This improvement in performance positions IC/DC as a new competitive learning-based method for addressing these less-studied CO problems, highlighting its potential.
@@
As shown in Tab. 1 and Tab. 2, simple implementations of diffusion models, whether trained with supervised learning [26] or reinforcement learning [5], struggle to generate high quality solutions without the aid of problem-specific search processes, which require significant expertise in the specific CO problem being addressed. This highlights the critical importance of our feasibility-enforced generation process, which integrates the flexibility of autoregressive methods into heatmap-based methods, allowing for the straightforward imposition of various constraints. In this regard, the proposed IC/DC approach significantly broadens the applicability of heatmap-based methods to a wide range of less-explored CO problems with diverse constraints.
5.3 Generalization capability
To assess the ability to generalize of IC/DC, we train our model on the ATSP-20 setting and evaluate it on ATSP instances with different numbers of nodes. Left side of Fig. 3 shows the result of this evaluation, comparing IC/DC with other learning-based methods. IC/DC demonstrates superior generalization capabilities, achieving $0\%$ optimality gap for 10 nodes. As the number of nodes increases, IC/DC maintains a lower optimality gap, with a notable $3.733\%$ gap at 30 nodes and relatively low gaps at higher node counts compared to other methods. These results suggests that IC/DC can effectively adapt to a range of different problems, confirming its ability to generalize across diverse, unseen distributions. (See Sec. E.1 for full results)
5.4 Reducing sampling steps for faster inference
Compared to autoregressive methods, diffusion models exhibit slower inference time due to the expensive reverse process. The parameterization in Eq. 7 enables inference with reduced sampling steps with a cost of accuracy through $p_{\theta}(X_{t-s}|X_{t},c)=\sum_{X_{0}}q(X_{t-s},X_{t}|X_{0})p_{\theta}(X_{0}|X_{t},c)$ [2], setting $t>1$. We examine the trade-offs associated with reduced sampling steps in the PMSP setting with 3 machines and 20 jobs. As shown in center and right of Fig. 3, the inference time decreases significantly as the number of sampling steps during inference $T_{I}$ decreases. IC/DC attains a $0\%$ optimality gap for both $T_{I}=\frac{1}{2}T$ and $T_{I}=\frac{1}{3}T$ settings with competitive inference speed. The trade-off reported here can be improved using a more sophisticated distillation algorithms [25]. (See Sec. E.2 for more results)
@@ 6 Conclusion
We propose a diffusion model framework, IC/DC, to tackle the CO problems with two distinct sets of items, showing competitive performance in both PMSP and ATSP. Our algorithm outperforms existing NCO baselines and generates feasible solutions that are difficult for heatmap-based methods. We believe the proposed IC/DC framework is flexible, and can be broadly applied to various domains; e.g., in Sec. D.4, we show its capability to address real-world CO problems with complex item and relationship features.
Despite its strong performance, IC/DC has a notable limitation: the GAT-based encoder we use requires $O(\max(|\mathcal{A}|,|\mathcal{B}|)^{2})$ memory complexity, demanding significant resources for large instances. Unfortunately, due to limited computational resources, we were not able to assess the performance of IC/DC on larger problem instances, which prevented us from fully exploring IC/DC’s scalability potential. To address these challenges, we plan to explore memory-efficient techniques in future work, such as those introduced by Zhu et al. [33].
Supplementary Material
@@ A.1 Joint Variational Upper Bound with EUBO
We derive the upper bound using the Evidence Upper Bound (EUBO) [14]
@@
$D_{KL}(\tilde{q}(X|c)\|p_{\theta}(X|c))$ $=\int\tilde{q}(X|c)\log\frac{\tilde{q}(X|c)}{p_{\theta}(X|c)}dX$ (23)
@@ A.2 KL divergence to Reward-weighted Diffusion Loss
We derive the reward-weighted diffusion loss from the objective in CLONING step. We aim to minimize the Kullback-Leibler (KL) divergence $D_{KL}(\tilde{q}(X|c)\parallel p_{\theta}(X|c))$ between the surrogate target and the distribution $p$ parameterized by $\theta$. however since evaluating the log-likelihood of the generative model $p_{\theta}(X|c)$ is difficult [16], we instead utilize a joint variational upper bound (see Sec. A.1):
$\mathrm{D}_{\mathrm{KL}}(\tilde{q}(X|c)\parallel p_{\theta}(X|c))\leq\mathrm{D}_{\mathrm{KL}}(\tilde{q}(X,Z|c)\parallel p_{\theta}(X,Z|c))$ (27)
@@
Therefore, we focus on minimizing the right-hand side of eq (27). By applying the diffusion process where $Z=X_{1:T}$, $X=X_{0}$, and $T$ is the diffusion step, the forward process is defined by $q(X_{1:T}|X_{0})$ with $X_{0}\sim\tilde{q}(X_{0}|c)$. We minimize the Reward-weighted Diffusion Loss $\mathcal{L}_{\text{RWD}}$:
$\mathcal{L}_{\text{RWD}}$ $=D_{KL}(\tilde{q}(X_{0:T}|c)\parallel p_{\theta}(X_{0:T}|c))$ (28)
$=D_{KL}(\tilde{q}(X_{0},X_{1:T}|c)\parallel p_{\theta}(X_{0:T}|c))$ (29)
$=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{1:T}\sim q(X_{1:T}|X_{0})}\left[\log\frac{q(X_{1:T}|X_{0})\tilde{q}(X_{0}|c)}{p_{\theta}(X_{0:T}|c)}\right]$ (30)
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[\log\frac{q(X_{1:T}|X_{0})}{p_{\theta}(X_{0:T}|c)}+\log\tilde{q}(X_{0}|c)\right]$ (31)
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[\log\frac{\prod_{t=1}^{T}q(X_{t}|X_{t-1})}{p_{\theta}(X_{T}|c)\prod_{t=1}^{T}p_{\theta}(X_{t-1}|X_{t},c)}+\log\tilde{q}(X_{0}|c)\right]$ (32)
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[-\log p_{\theta}(X_{T}|c)+\sum_{t=1}^{T}\log\left(\frac{q(X_{t}|X_{t-1})}{p_{\theta}(X_{t-1}|X_{t},c)}\right)+\log\tilde{q}(X_{0}|c)\right]$ (33)
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[-\log p_{\theta}(X_{T}|c)+\sum_{t=2}^{T}\log\left(\frac{q(X_{t}|X_{t-1})}{p_{\theta}(X_{t-1}|X_{t},c)}\right)+\log\left(\frac{q(X_{1}|X_{0})}{p_{\theta}(X_{0}|X_{1},c)}\right)+\log\tilde{q}(X_{0}|c)\right]$ (34)
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\Bigg{[}-\log p_{\theta}(X_{T}|c)$ (35)
$\quad\quad\quad\quad\quad\quad\quad\ \ +\sum_{t=2}^{T}\log\left(\frac{q(X_{t-1}|X_{t},X_{0})}{p_{\theta}(X_{t-1}|X_{t},c)}\cdot\frac{q(X_{t}|X_{0})}{q(X_{t-1}|X_{0})}\right)+\log\left(\frac{q(X_{1}|X_{0})}{p_{\theta}(X_{0}|X_{1},c)}\right)+\log\tilde{q}(X_{0}|c)\Bigg{]}$ (36)
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\left[\log\frac{q(X_{T}|X_{0})}{p_{\theta}(X_{T}|c)}+\sum_{t=2}^{T}\log\left(\frac{q(X_{t-1}|X_{t},X_{0})}{p_{\theta}(X_{t-1}|X_{t},c)}\right)-\log p_{\theta}(X_{0}|X_{1},c)+\log\tilde{q}(X_{0}|c)\right]$ (37)
$=\mathbb{E}_{X_{0}\sim\tilde{q},X_{1:T}\sim q}\Bigg{[}\underbrace{D_{KL}(q(X_{T}|X_{0})\|p_{\theta}(X_{T}|c))}_{L_{T}}+\sum_{t=2}^{T}\underbrace{D_{KL}(q(X_{t-1}|X_{t},X_{0})\|p_{\theta}(X_{t-1}|X_{t},c))}_{L_{t-1}}$ (38)
$\quad\quad\quad\quad\quad\quad\quad\ \ +\underbrace{D_{KL}(\tilde{q}(X_{0}|c)\|p_{\theta}(X_{0}|X_{1},c))}_{L_{0}}\Bigg{]}$ (39)
If $T$ is sufficiently large, $L_{T}$ will approach zero [2]. Also, $L_{0}$ can be derived as CE loss:
$L_{0}$ $=\sum\tilde{q}(X_{0}|c)\log\frac{\tilde{q}(X_{0}|c)}{p_{\theta}(X_{0}|X_{1},c)}=\sum\tilde{q}(X_{0}|c)\log\tilde{q}(X_{0}|c)-\log p_{\theta}(X_{0}|X_{1},c))$ (40)
$=CE(\tilde{q},p_{\theta})-H(\tilde{q})\approx\mathcal{L}_{\text{prd}}-H(\tilde{q})$ (41)
@@
$\therefore\mathcal{L}_{\text{RWD}}\approx\mathcal{L}_{\text{VB}}+\mathcal{L}_{\text{prd}}$ (42)